a. With Laplace transformers solve x"+4x'+20x=0 ; x(0)=4 & x'(0)=-5

b. With Laplace transformers solve x'=2x+2y and y'=2x-y ; x(0)=1 & y(0)=2
Answer #1

a) 4x! + 20x=0 x 10) = X(o)= -5 Laplace on both side 16x7 + 4.L{x} + 20 24x} =0 $?? [x] - SX10) - x?(0) +4 [Slaxy - XI0)) +S+5 22x3- 3+5 5²-5-6 S- (5-12-23 24x3 = X(t)= ? Cossht insht & 11 x 3 en x(t) = eth lossnt + Letrasins, ht (5+1) 1fy) - , 149